1. Show that the rate of change of the total angular momentum of a system of particles is equal to the resultant torque exerted by all external forces which act on the system. 
1
Expert's answer
2021-01-11T11:41:22-0500

Let the particle of mass m, is moving with velocity "\overrightarrow{v}" then linear momentum of the particle will be "\overrightarrow{p}" .and particle is at a distance "\overrightarrow{r}" from the origin.

Let the angular momentum of the particle is "\overrightarrow{L}."



So, angular momentum of the particle will be "\overrightarrow{L}=\overrightarrow{r}\times \overrightarrow{p}"

"\Rightarrow \overrightarrow{L}=|\overrightarrow{r}||\overrightarrow{p}|\sin\theta \hat{n}"

where "\theta" is the angle between "\overrightarrow{r}" and "\overrightarrow{p}" and "\hat{n}" is the unit vector perpendicular to "\overrightarrow{r}" and "\overrightarrow{p}"

Now, taking the time derivative of "\overrightarrow{L}=\overrightarrow{r}\times \overrightarrow{p}"

"\Rightarrow \frac{d\overrightarrow{L}}{dt}=\frac{d \overrightarrow{r}}{dt}\times \overrightarrow{p}+\overrightarrow{r}\times \frac{{d\overrightarrow{p}}}{dt}"

from newton's second law of motion,"\Sigma \overrightarrow{F}=\frac{d\overrightarrow{p}}{dt}"

Hence, "\frac{d\overrightarrow{L}}{dt}=\overrightarrow{r}\times \overrightarrow{F}"

But we know that "\overrightarrow{\tau}= \overrightarrow{r}\times \overrightarrow{F}"

Hence, "\overrightarrow{\tau}=\frac{d\overrightarrow{L}}{dt}"